### v1

# Resolution of Simpson’s paradox via the common cause principle

Abstract
Simpson’s paradox is an obstacle to establishing a probabilistic association between two events $a_{1}$ and $a_{2}$, given the third (lurking) random variable $B$. We focus on scenarios when the random variables $A$ (which combines $a_{1}$, $a_{2}$, and their complements) and $B$ have a common cause $C$ that need not be observed. Alternatively, we can assume that $C$ screens out $A$ from $B$. For such cases, the correct association between $a_{1}$ and $a_{2}$ is to be defined via conditioning over $C$. This set-up generalizes the original Simpson’s paradox. Now its two contradicting options simply refer to two particular and different causes $C$. We show that if $B$ and $C$ are binary and $A$ is quaternary (the minimal and the most widespread situation for valid Simpson’s paradox), the conditioning over any binary common cause $C$ establishes the same direction of the association between $a_{1}$ and $a_{2}$ as the conditioning over $B$ in the original formulation of the paradox. Thus, for the minimal common cause, one should choose the option of Simpson’s paradox that assumes conditioning over $B$ and not its marginalization. For tertiary (unobserved) common causes $C$ all three options of Simpson’s paradox become possible (i.e. marginalized, conditional, and none of them), and one needs prior information on $C$ to choose the right option.
I Introduction
Simpson’s paradox was discovered more than a century ago pearson ; yule , generated a vast literature, and is well-recognized in several fields cohen ; simpson ; blyth ; lindley ; barigelli ; barigelli2 ; pearl ; pearl2 ; hernan ; appleton ; armi ; agresti ; cohen_uncertainty . This counter-intuitive effect is said to limit one’s ability to draw obvious conclusions from probabilistic data. The effect is important because it demands more than simply extracting relative frequencies from data; e.g. it necessitates looking at exchangeability or causality lindley .
To formulate the paradox in its simplest form, assume three binary random variables $A_{1}=\{a_{1},\bar{a}_{1}\}$, $A_{2}=\{a_{2},\bar{a}_{2}\}$, $B=\{b,\bar{b}\}$. The target event is $a_{1}$, and we would like to know how it is influenced by $A_{2}$ which occurs at an earlier time than the time of $A_{1}$: $t_{A_{2}}\leq t_{A_{1}}$. This can be done by looking at conditional probability. For
$p(a_{1}|a_{2})<p(a_{1}|\bar{a}_{2}),$ (1)
which is equivalent to $p(a_{1})<p(a_{1}|\bar{a}_{2})$, we would conclude that $\bar{a}_{2}$ enables $a_{1}$. However, (1) is compatible with
$p(a_{1}|a_{2},b)$ $>p(a_{1}|\bar{a}_{2},b),$ (2)
$p(a_{1}|a_{2},\bar{b})$ $>p(a_{1}|\bar{a}_{2},\bar{b}),$ (3)
where $B$ also occured in an earlier time: $t_{B}\leq t_{A_{1}}$. For example, $a_{1}$ and $\bar{a}_{1}$ can be positive and negative results, while $a_{2}$ and $\bar{a}_{2}$ mean two different medical treatments; $b$ and $\bar{b}$ can then refer to different features of patients, e.g. their sex ($b={\rm male}$, $\bar{b}={\rm female}$) or age ($b={\rm old}$, $\bar{b}={\rm young}$) lindley ; see Appendix A for more examples. Now since (2, 3) hold for each value of $B$ we should perhaps conclude that ${a}_{2}$ enables $a_{1}$ in contrast to (1). Decision-makers would not know whether to apply (1) or (2, 3). This is Simpson’s paradox. For the paradox, $A_{2}$ and $B$ must be dependent variables  [^1] We leave aside the following pertinent problem; see cohen_uncertainty for details. If probabilities are extracted from finite populations, the more conditioned version (2, 3) is less reliable, because it is extracted from a smaller population. For us all probability-providing populations will be sufficiently large. .
Over time, several resolutions to the paradox have been proposed. Barigelli and Scozzafava barigelli ; barigelli2 proposed to replace (1) by $p(a_{2})p(a_{1}|a_{2})<p(\bar{a}_{2})p({a}_{1}|\bar{a}_{2})$, i.e. to interchanging $A_{1}$ and $A_{2}$ in (1). Then it is easy to see that its inversion under additional conditioning over $B$ is impossible. While (1) stands for prediction, i.e. aiming at $a_{2}$ (and not at $\bar{a}_{2}$) will more likely produce $a_{1}$ (than $\bar{a}_{1}$), the proposal by Ref. barigelli ; barigelli2 looks for retrodiction. Though retrodicting (in contrast to predicting) does not suffer from Simpson’s paradox, retrodicting and predicting are different things, hold different intuitions, and cannot generally be substituted for each other. In a related proposal, Rudas also sought to change the criterion (1) so that it does not allow inversion after additional conditioning over $B$, but still has several reasonable features rudas . The proposal is to employ $p(a_{2})[p(a_{1}|a_{2})-p(\bar{a}_{1}|a_{2})]<p(\bar{a}_{2})[p(a_{1}|\bar{a}_{2})-p(\bar{a}_{1}|\bar{a}_{2})]$ instead of (1) rudas . Notice the conceptual relation of this with the previous proposal  [^2] An unnatural point of both these proposals is that they depend on the ratio $p(a_{2})/p(\bar{a}_{2})$; e.g. for the example mentioned after (3) this means that the treatment that was applied more has better chances to be accepted. This drawback is also acknowledged in rudas . .
According to Lindley and Novick, the paradox may be resolved by going beyond probabilistic considerations (as we do here as well) and by employing the notion of exchangeability or causality lindley . Within that proposal, the data generally provides only propensities, and one needs additional assumptions of sample homogeneity (exchangeability) for equating propensities with probabilities even for a large sample size. They studied two examples [see Examples 1 and 2 in Appendix A], where natural assumptions on exchangeability lead to preferring (1) for Example 2 and (2, 3) for Example 1 lindley . Lindley and Novick also proposed that the same results can be found via using causality instead of exchangeability lindley . Pearl elaborated this proposal showing that the examples of Ref. lindley can be represented via time-ordered direct acyclic graphs (TODAG) pearl ; pearl2 ; see Fig. 1 and Appendix A for details. The suggestion is to condition over $B$ [hence using (2, 3)] if $B$ causes both $A_{1}$ and $A_{2}$. This is because conditioning over the cause reduces spurious correlations. In contrast, it is advised to use (1) if $B$ is an effect of $A_{2}$, but still a cause of $A_{1}$ pearl ; pearl2  [^3] The intuition of this suggestion is seen in the extreme case when $B$ screens $A_{1}$ and $A_{2}$ from each other, i.e. $A_{1}$, $B$ and $A_{2}$ form a Markov chain. Then the conditional probability $p(A_{1}|A_{2},B)=p(A_{1}|B)$ will not depend on $A_{2}$ begging the original question in (1). This reasoning was generalized as the back-door criterion pearl . . For the two examples considered in lindley , Refs. pearl ; pearl2 make similar recommendations. The basis of these recommendations was criticized in armi .
Note, however, that the examples presented in Ref. lindley (and reproduced in Appendix A) do not need to support any TODAG. The data of these examples support an alternative interpretation that $B$ and $A=(A_{1},A_{2})$ do not have any causal influence on each other. There are also more examples of this type; see Appendix A. This is the starting point of our analysis: once $B$ and $A=(A_{1},A_{2})$ do not influence each other casually, it is natural to assume—following the common cause principle reich ; suppes —that there is an unobserved random variable $C$, which is a common cause for $A$ and $B$; see section II. An alternative assumption that leads to the same results is that $C$ causes only $A$, but it also screens out $A$ from $B$.
The introduction of this principle allows us to reformulate and extend Simpson’s paradox: (1) and (2, 3) —along with many other options—are particular cases that refer to particular choices of the unobserved cause $C$; see the next section. Now the paradox seems to be further from resolution than ever before. However, when $A_{1}$, $A_{2}$, $B$ and $C$ are binary (the minimal set-up of the paradox), and provided that the paradox holds (i.e. (1) and (2, 3) hold), all causes $C$ imply that the decision making is to be taken according to (2, 3); see the next section. The last section summarizes this work and discusses tertiary causes.
Figure 1: Directed acyclic graphs between random variables $A=(A_{1},A_{2})$, $B$ and $C$ involved in discussing Simpson’s paradox. The first and second graphs were studied in Refs. pearl ; pearl2 . They are discussed in section I and Appendix A. The third or fourth graphs are basic assumptions of this work; see (4). In the first graph, $B$ influences $A_{1}$ and $A_{2}$, but $B$ is not the common cause in the strict sense, because there is an influence from $A_{2}$ to $A_{1}$. A similar interpretation applies to the second graph. We emphasize that the joint probability $p(A_{1},A_{2},B)$ for the first and second graphs has the same form, i.e. such graphs are extra constructions involved for an additional interpretation of data. In contrast, the third and fourth graph imply a definite (but the same for both graphs) limitation on the joint probability $p(A_{1},A_{2},B,C)$, which is expressed by (4).
II Common cause principle and reformulation of Simpson’s paradox
The common cause for $A=(A_{1},A_{2})$ and $B$ means reich ; suppes
$p(A_{1},A_{2},B|C)=p(A_{1},A_{2}|C)p(B|C),$ (4)
where this screening relation holds for all values assumed by $A_{1}$, $A_{2}$, $B$ and $C$. The times of these variables naturally hold $t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}},t_{B}]$. The same (4) applies if $C$ causes $A$ and screens $A$ from $B$. Now $t_{B}<t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}}]$. These two scenarios are shown in Fig. 1 as (resp.) the third and fourth graphs. Appendix A provides several examples for a causing (or screening) variable $C$ in the context of Simpson’s paradox.
Note from (4) that $C$ gets rid of the conditional dependence on $B$ in $p(A_{1},A_{2}|B,C)$. Thus, the correct way of looking at the association between $a_{1}$ and $a_{2}$ is to check the sign of $p(a_{1}|a_{2},C)-p(a_{1}|\bar{a}_{2},C)$ for all values of $C$. To support this point, we note that (4) has an important implication in the context of (1). (This implication generalizes the argument given in suppes .) Assume that $p(a_{2},b,c)>0$ for all values $c$ of $C$. Note from (4) that there exists an event $c$ such that $p(a_{1}|a_{2},b)\leq p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$, and an event $c^{\prime}$ such that $p(a_{1}|a_{2},b)\geq p(a_{1}|a_{2},b,c^{\prime})=p(a_{1}|a_{2},c^{\prime})$. Hence, if conditioning over $b$ facilitates (hinders) the association between $a_{1}$ and $a_{2}$, then conditioning over $c$ ($c^{\prime}$) is not worse in this facilitation (hindering)  [^4] To deduce the first relation assume that $p(a_{1}|a_{2},b)<p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$ for all $c$, multiply both parts by $p(a_{2},b,c)>0$, sum over $c$ and get contradiction $p(a_{1},a_{2},b)<p(a_{1},a_{2},b)$. Likewise for the second relation..
After the above reformulation, Simpson’s paradox seems even less resolvable since $C$ is not observed. Indeed, there are common causes that reproduce (1), those that reproduce (2, 3), but there are many other possibilities. Common causes that are close to $B$ ($C\approx B$) imply option (2, 3) of the paradox, while $C\approx A$ leads to (1). These conclusions are based on the fact that (4) holds exactly for $C=B$ and $C=A$. Thus, Simpson’s paradox is not a choice between two options (2, 3) and (1), it is a choice between many options given by different common causes $C$.
The following theorem shows a definite statement for all binary causes. The message of the theorem is that once we know that $C$ is binary, then the correct decision is (2, 3).
Theorem: If $A_{1}$, $A_{2}$, $B$ and $C=\{c,\bar{c}\}$ are binary, and provided that (1) and (2, 3) are valid, all causes $C$ hold
$p(a_{1}|a_{2},c)>p(a_{1}|\bar{a}_{2},c),\qquad p(a_{1}|a_{2},\bar{c})>p(a_{1}|\bar{a}_{2},\bar{c}),$ (5)
i.e. all $C$ holding (4) predict the same sign of association between $a_{1}$ and $a_{2}$ as (2, 3). The main idea of proving (5) is inverting (4), i.e. presenting it as
$p(a_{1}|a_{2},c)$ $=$ $\frac{p(\bar{b}|\bar{c})p(a_{1}|a_{2},b)p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2})}{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}$ (6)
$=$ $p(a_{1}|a_{2},\bar{b})+\frac{p(\bar{b}|\bar{c})p(b|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]}{1-p(\bar{b}|\bar{c})-p(b|a_{2})},$ (7)
$p(c|a_{2})$ $=$ $\frac{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}{p(b|c)+p(\bar{b}|\bar{c})-1}=\frac{p(\bar{b}|\bar{c})+p(b|a_{2})-1}{p(b|c)+p(\bar{b}|\bar{c})-1}$ (8)
where unknown quantities $p(a_{1}|a_{2},c)$ and $p(c|a_{2})$ are represented via known ones (i.e. $p(A_{1},A_{2},B)$) and free parameters $p(B|C)$. Eqs. (7, 8) hold upon changing $a_{2}$ by $\bar{a}_{2}$. Eqs. (7, 8) are deduced in Appendix B via specific notations that should be useful when dealing with (4) more generally, i.e. for non-binary $C$.
The rest of the proof is algebraic but non-trivial. It also works out and employs constraints (13, 18) on Simpson’s paradox itself. Expanding both sides of (1),
$p(a_{1}|a_{2})=p(a_{1}|a_{2},b)p(b|a_{2})+p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2}),$ (9)
$p(a_{1}|\bar{a}_{2})=p(a_{1}|\bar{a}_{2},b)p(b|\bar{a}_{2})+p(a_{1}|\bar{a}_{2},\bar{b})p(\bar{b}|\bar{a}_{2}),$ (10)
and using there (2, 3) we subtract both sides of (1) from each other and find:
$p(a_{1}|a_{2},b)+p(\bar{b}|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]<p(a_{1}|\bar{a}_{2},b)+p(\bar{b}|\bar{a}_{2})[p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)].$ (11)
We return to (2, 3) and note that we can assume without loosing generality
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|a_{2},b).$ (12)
Eqs. (9, 10) imply that for the validity of (13, 12) it is necessary to have $p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)$, which together with (2, 3, 12) can be written as
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)>p(a_{1}|\bar{a}_{2},{b}).$ (13)
Now (1, 9, 10) can be written as
$p(a_{1}|a_{2},b)+[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<p(a_{1}|\bar{a}_{2},b)+[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (14)
$p(a_{1}|a_{2},\bar{b})-p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<p(a_{1}|\bar{a}_{2},\bar{b})-p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (15)
where (14) and (15) are equivalent. Now (14) and (15) imply from (13) (resp.)
$[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (16)
$p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))>p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)).$ (17)
As checked directly, Eqs. (16, 17) lead to
$p(b|a_{2})>p(b|\bar{a}_{2}).$ (18)
Now we return to (8) and assume there ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$, which leads to ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}<0$ from (8). Writing down from (8) the formula for $p(c|\bar{a}_{2})$ and making the same assumption we get ${p(\bar{b}|\bar{c})+p(b|\bar{a}_{2})-1}<0$. Now look at (7) and its analog obtained via $a_{2}\to\bar{a}_{2}$, and use there these two results together with (17, 18) and (13) to deduce the first inequality in (5) under assumption ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$. It should be obvious that the second inequality in (5) holds under the same assumption since we nowhere used any specific feature of $c$ compared to $\bar{c}$.
For ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}>0$ we need to use instead of (7) another form of (6)
$p(a_{1}|a_{2},c)=p(a_{1}|a_{2},b)-\frac{[1-p(\bar{b}|\bar{c})]p(\bar{b}|a_{2})[p(a_{1}|a_{2},b)-p(a_{1}|a_{2},\bar{b})]}{p(\bar{b}|\bar{c})+p(b|a_{2})-1}.$ (19)
The rest is similar to the above: we proceed via (16, 18) and (13) and deduce (5) from (8), (19) and the analog of (19) obtained via $a_{1}\to\bar{a}_{2}$.
III Summary
We addressed Simpson’s paradox: the problem of setting up an association between two events $a_{1}$, $a_{2}$ given the lurking variable $B$. The paradox provides two plausible but opposite conclusions for the same situation; see (1) and (2, 3). Thus, there are three possible outcomes: either the first option is correct, the second option is correct, or none of them are correct.
We focus on cases, when there is a common cause $C$ for $B$ and $A=(A_{1},A_{2})$ (which combines $a_{1}$, $a_{2}$ and their complements). Alternatively, $C$ screens out $A$ from $B$. These cases include those in which there is no causal influence from $A$ to $B$, as well as from $B$ to $A$. Hence, correlations between $A$ and $B$ are to be explained via the common cause $C$, which is a statement of the common cause principle reich ; suppes . Now the association between $a_{1}$ and $a_{2}$ is to be decided by looking at $p(a_{1}|a_{2},c)$ for various values of $C$. This task is normally difficult given the fact that $C$ is frequently not fully known and is not observed. However, provided that $A_{1}$, $A_{2}$, $B$ and $C$ are binary, $p(a_{1}|a_{2},c)$ shows the same association as the option (2, 3) of Simpson’s paradox. In this sense, Simpson’s paradox is resolved in the binary situation, provided that the situation allows a binary cause or a binary screening variable.
Starting from a tertiary $C$, all three options of Simpson’s paradox become possible: there are common causes $C$ that support (1), those which support (2, 3), and eventually (random) cause variables $C={c_{1},c_{2},c_{3}}$ for which $p(a_{1}|a_{2},c_{i})-p(a_{1}|\bar{a}_{2},c_{i})$ has different signs for different values of $i=1,2,3$. Our numerical examples showing these possibilities are contained in Supplementnumerics . Hence, already for the tertiary cause one needs prior information on the common cause to decide on the solution of Simpson’s paradox. Alternatively, we can infer this unknown cause via e.g. one of methods proposed recently koc ; hovh2023 . It is not excluded that such inference methods will provide further information on the solution of Simpson’s paradox. We hope to discuss this problem elsewhere.
Acknowledgements.
We were supported by HESC of Armenia, grants 24FP-1F030 and 21AG-1C038.
Appendix A Examples of Simpson’s paradox
We collected several examples of the paradox that are scattered in the literature. We discuss them employing our notations in (13) emphasizing (whenever relevant) the existence of the common cause (or screening) variable $C$.
Example 1. Medical treatment lindley . $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the recovery rate of medical patients: $a_{1}={\rm recovery}$, $\bar{a}_{1}={\rm no~{}recovery}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to a specific medical treatment: $a_{2}={\rm treatment}$, $\bar{a}_{2}={\rm no~{}treatment}$. $B=\{b,\bar{b}\}$ is the sex of patients: $b={\rm male}$, $\bar{b}={\rm female}$. The times to which the random variables $A_{1}$, $A_{2}$ and $B$ refer clearly hold $t_{B}<t_{A_{2}}<t_{A_{1}}$. Ref. lindley argued for this example that a better method to decide on the association between $A_{1}$ and $A_{2}$ is (2, 3).
Given that this example holds Simpson’s paradox and assuming that it can be described via a time-ordered direct acyclic graph (TODAG), there is only one TODAG that accounts for it: $B\to A_{2}\to A_{1}\leftarrow B$; see the first graph in Fig. 1. Then $B$ causes both $A_{1}$ and $A_{2}$, and—according to lindley ; pearl ; pearl2 —it is natural to condition over it leading to the decision rule (2, 3).
It is possible as well that this situation is not described by any TODAG, because there is no direct causal influence between $A=(A_{1},A_{2})$ and $B$. Then one possible common cause for $A$ and $B$ is the genetic makeup.
Example 2. Plant yield lindley . $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the yield of a single plant: $a_{1}={\rm high}$, $\bar{a}_{1}={\rm low}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to the variety (color) of the plant: $a_{2}={\rm dark}$, $\bar{a}_{2}={\rm light}$. $B=\{b,\bar{b}\}$ refers to the height of the plant: $b_{1}={\rm tall}$, $\bar{b}={\rm low}$. The times hold $t_{A_{2}}<t_{B}<t_{A_{1}}$. Given that this situation holds Simpson’s paradox and assuming that it can be described via a time-ordered direct acyclic graph (TODAG), there is only one TODAG that accounts for it: $A_{2}\to B\to A_{1}\leftarrow A_{2}$; see the second graph in Fig. 1. Now since $B$ blocks one possible path from $A_{2}$ to $A_{1}$, Ref. pearl2 proposes not to condition over it suggesting that the correct decision rule is (1). This also coincides with the suggestion made in Ref. lindley concerning this example.
Again, the data itself does not assume any direct causal influence between $A=(A_{1},A_{2})$ and $B$, i.e. this example need not be described by any TODAG. One possible common cause $C$ for $A$ and $B$ is the genetic makeup.
Example 3. This example is taken from Ref. appleton . Here $A_{1}$ amounts to the survival of a person (i.e. reaching a certain age), the binary $A_{2}$ refers to smoker versus non-smoker, while $B$ means the age group. It turns out that the aggregated data (1) shows that smoking is beneficial for survival. This conclusion is reversed, once the age group $B$ is introduced, i.e. (2, 3) is employed instead of (1). The peculiar point of this example is that according to Ref. appleton , $B$ (i.e., the number of people of a certain age) was influenced by $A_{2}$. Hence this example is of the same type as Example 2 and could be described by the TODAG $A_{2}\to B\to A_{1}\leftarrow A_{2}$. However, marginalizing over $B$ (i.e. employing (1)) will lead to a wrong interpretation that smoking is good for survival.
Example 4. Snow tires provide cars with better traction in snowy and icy road conditions. However, nationally in the US, cars fitted with snow tires are more likely to have accidents in snowy and icy conditions norton . $A_{1}=\{a_{1}={\rm accident},\bar{a}_{1}={\rm no-accident}\}$, $A_{2}=\{a_{2}={\rm changed~{}tires},\bar{a}_{2}={\rm not~{}changed}\}$, $B=\{{\rm states}\}$. Here the choice of the state (warm or cold) has a direct causal link to accidents in winter conditions. Now snow tires tend to be fitted to cars only in snowy winter months and in states with colder weather. Cars in warmer months and in states with warmer weather are much less likely to have accidents in snowy and icy conditions. Plausibly, there is a random variable, $C=\{{\rm good~{}weather~{}conditions},{\rm bad~{}weather~{}conditions}\}$, which causes $A$, and screens $A$ from $B$: $p(A|CB)=p(A|C)$. The times are distributed as $t_{B}<t_{C}<t_{A_{2}}<t_{A_{1}}$.
Example 5. We need to choose between two hospitals 1 and 2: $A_{1}=\{a_{1}={\rm recovered},\bar{a}_{1}={\rm not~{}recovered}\}$, $A_{2}=\{a_{2}={\rm hospital~{}1},\bar{a}_{2}={\rm hospital~{}2}\}$, $B=\{{\rm first~{}half-year},{\rm second~{}half-year}\}$, $C=\{{\rm types~{}of~{}illness}\}$. Here we do not expect direct causal influence from $B$ to $A$, if (as we assume) the hospitals do not treat seasonal illnesses. We expect that $C$ causes $A$, and screens it from $B$.
Note that the data from which the probabilities for Simpson’s paradox are calculated is the number of patients $N(A_{1},A_{2},B)$ that came to the hospital. Simpson’s paradox does not exist—moreover, the decisions based on conditioning and on marginalization of $B$ are equivalent—if within each season the hospitals accept an equal number of patients: $\sum_{A_{1}}N(A_{1},A_{2},B)=\sum_{A_{1}}N(A_{1},A_{2},B)$ for each value of $B$. This creates a conceptual possibility for judging between the hospitals. This is however not realistic, because imposing on these hospitals an equal number of patients can disturb their usual (normal) functioning.
Example 6. Simpson’s paradox is realized when comparing scores of professional athletes, e.g. the batting averages of baseball players ross . Here $A_{1}$ refers to a score of an athlete in a game, e.g. $A_{1}={\rm high~{}score,low~{}score}$, $A_{2}$ denotes concrete athletes, while $B$ is the time-period (e.g. playing season). The causing variable $C$ can refer to the psychological and physical state of an athlete that influences his/her game success, and the number of games he/she participated in each season.
Example 7. Here the COVID-19 death rates are compared in Italy and China stanford . According to the data, aggregated death rates in Italy are higher than in China, but in each age group, the death rates are higher in China. Specific details of age groups were not reported in Ref. stanford ; for this example, we take it binary: above 65 years and below it. We assign the death rate – $A_{1}$, country – $A_{2}$, age group – $B$. We propose that the healthcare system can be the common cause $C$ here: a more affordable healthcare system may cause a higher proportion of older people in the country’s society. Indeed, for 2019, Italy had a larger percentage of people aged above 65 than China: 24.05 % versus 12.06 %. On the other hand, the healthcare system will influence death rates in all age groups. When faced with the COVID-19 pandemic, older people have a higher death rate. If the proportion of older people is high, this may result in higher overall death rates, though within each age group, the death rate difference is the opposite.
Appendix B Matrix notations for inverting the common cause equation
Here we develop matrix notations for inverting the common cause equation:
$p(A_{1},A_{2},B)=\sum_{C}p(A_{1},A_{2},C)p(B|C),$ (20)
where the summation goes over all values of $C$. We work for the case when the variables $A_{1}$, $A_{2}$, $B$ and $C$ are binary, though the matrix notations we introduce below are useful more generally.
Eq. (20) can be written in matrix form
$\begin{pmatrix}[ik1]\\ [ik2]\end{pmatrix}=\begin{pmatrix}(1|1)&(1|2)\\ (2|1)&(2|2)\end{pmatrix}\begin{pmatrix}(ik1)\\ (ik2)\end{pmatrix}.$ (27)
where $ik=11,12,21,22$ and the following notations were introduced
$\begin{split}&[111]\equiv p(a_{1},a_{2},b),\hskip 5.0pt[121]\equiv p(a_{1},\bar{a}_{2},b),\dots,\\ &(111)\equiv p(a_{1},a_{2},c),\hskip 5.0pt(121)\equiv p(a_{1},\bar{a}_{2},c),\dots,\\ &(1|1)\equiv p(b|c),\hskip 5.0pt(2|1)\equiv p(\bar{b}|c),\dots,\\ &D=(2|2)+(1|1)-1.\end{split}$ (28)
Inversion of the Eq. (27) gives
$\begin{pmatrix}(ik1)\\ (ik2)\end{pmatrix}=\frac{1}{D}\begin{pmatrix}(2|2)&-(1|2)\\ -(2|1)&(1|1)\end{pmatrix}\begin{pmatrix}[ik1]\\ [ik2]\end{pmatrix}.$ (35)
Eq. (35) implies
$\begin{pmatrix}(i|k1)\{1|k\}\\ (i|k2)\{2|k\}\end{pmatrix}=\frac{1}{D}\begin{pmatrix}(2|2)&-(1|2)\\ -(2|1)&(1|1)\end{pmatrix}\begin{pmatrix}[i|k1][1|k]\\ [i|k2][2|k]\end{pmatrix},$ (42)
where analogously to (28) we introduced the following notations:
$\begin{split}&[1|11]\equiv p(a_{1}|a_{2},b),\hskip 5.0pt[1|21]\equiv p(a_{1},\bar{a}_{2},b),\dots,\\ &(1|11)\equiv p(a_{1}|a_{2},c),\hskip 5.0pt(1|21)\equiv p(a_{1}|\bar{a}_{2},c),\dots,\\ &[1|1]\equiv p(b|a_{2}),\hskip 5.0pt[2|1]\equiv p(\bar{b}|a_{2}),\dots,\\ &\{1|1\}\equiv p(c|a_{2}),\hskip 5.0pt\{2|1\}\equiv p(\bar{c}|a_{2}),\dots.\end{split}$ (43)
The matrix relation (42) results in
$\begin{split}(i|k1)\{1|k\}=\frac{(2|2)}{D}[i|k1][1|k]+\frac{(2|2)-1}{D}[i|k2][2|k].\end{split}$ (44)
Using (44, 35) we get relations employed in the main text:
$(i|k1)$ $=\frac{(2|2)[i|k1][1|k]+((2|2)-1)[i|k2][2|k]}{(2|2)[1|k]+((2|2)-1)[2|k]},$ (45)
$\{1|k\}$ $=\frac{1}{D}(2|2)[1|k]+\frac{1}{D}((2|2))-1)[2|k],$ (46)
$(i|k2)$ $=\frac{((1|1)-1)[i|k1][1|k]+(1|1)[i|k2][2|k]}{((1|1)-1)[1|k]+(1|1)[2|k]},$ (47)
$\{2|k\}$ $=\frac{1}{D}((1|1)-1)[1|k]]+[1|1][2|k]).$ (48)