### v2

# Resolution of Simpson’s paradox via the common cause principle

Abstract
Simpson’s paradox is an obstacle to establishing a probabilistic association between two events $a_{1}$ and $a_{2}$, given the third (lurking) random variable $B$. We focus on scenarios when the random variables $A$ (which combines $a_{1}$, $a_{2}$, and their complements) and $B$ have a common cause $C$ that need not be observed. Alternatively, we can assume that $C$ screens out $A$ from $B$. For such cases, the correct association between $a_{1}$ and $a_{2}$ is to be defined via conditioning over $C$. This setup generalizes the original Simpson’s paradox: now its two contradicting options refer to two particular and different causes $C$. We show that if $B$ and $C$ are binary and $A$ is quaternary (the minimal and the most widespread situation for the Simpson’s paradox), the conditioning over any binary common cause $C$ establishes the same direction of association between $a_{1}$ and $a_{2}$ as the conditioning over $B$ in the original formulation of the paradox. Thus, for the minimal common cause, one should choose the option of Simpson’s paradox that assumes conditioning over $B$ and not its marginalization. The same conclusion is reached when Simpson’s paradox is formulated via 3 continuous Gaussian variables: within the minimal formulation of the paradox (3 scalar continuous variables $A_{1}$, $A_{2}$, and $B$), one should choose the option with the conditioning over $B$.
{IEEEImpStatement}
Simpson’s paradox is a genuine decision-making problem, and not just an artefact of a deficient data-gathering process. This also follows from our re-estimation of its frequency in an unbiased data generation process. Hence, the paradox can lead to incorrect recommendations in various fields ranging from artificial intelligence to psychology. This paper shows that the paradox can be resolved if the minimal (not necessarily observed) common cause (or screening variable) can be assumed for the involved random variables. The result holds both in discrete and continuous settings, which relates to linear regressions. On several non-trivial examples from the published literature (smoking, COVID-19, etc), we show that such an assumption is plausible and leads to new predictions. Given deep relations between the common cause principle and basic unsupervised learning methods (Non-negative Matrix factorization, Probabilistic Latent Dirichlet indexing), we expect that our results will facilitate connections between decision-making and machine learning algorithms.
{IEEEkeywords}
causality, confounding, decision making, probabilistic association
1 Introduction
[…]
Simpson’s paradox was discovered more than a century ago [1, 2], generated a vast literature, and is well-recognized in several fields including, statistics, epidemiology, psychology, social science, etc. [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. This counter-intuitive effect limits the ability to draw conclusions from probabilistic data. The effect is important because it demands more than simply extracting relative frequencies from data; e.g. it necessitates looking at exchangeability [9] or causality [7, 8, 9, 13, 14].
The paradox starts with two random variables $A$ and $B$. Now $A=(A_{1},A_{2})$ contains control variable $A_{2}$ and the target variable $A_{1}$, while $B$ is a side random variable that correlates with both $A_{1}$ and $A_{2}$. The meaning of $A$ and $B$ is clarified via examples presented below. If there is no information on the outcome of $B$, the behavior of $A$ can be studied on two levels. The first (aggregated) level is that of marginal probabilities $p(A=a)$. The second level is finer-grained and is represented by conditional probabilities $p(A=a|B=b)$ for all possible values of $B$. Simpson’s paradox amounts to certain relations between those probabilities; see section 2 for details. It states that no decision-making is possible, because conclusions drawn from probabilities on different levels contradict each other. Without Simpson’s paradox, decision-making can proceed at the aggregate level, because looking at the fine-grained level is either redundant or inconclusive. Thus, Simpson’s paradox first and foremost involves decision-making. Moreover, it demonstrates limitations of the sure-thing principle [5], a pillar of traditional decision making [25, 26, 27]. A recent review of the sure-thing principle (and its limitations other than Simpson’s paradox) can be found in Ref. [28]. Limitations of probabilistic decision-making are important for the modern artificial intelligence (probability models, uncertainty estimation, etc).
In section 2, Simpson’s paradox is defined in detail, and previous efforts to resolve it in several specific situations are reviewed and criticized. In particular, we show that while certain previous solutions of the paradox assumed the existence of (causally-sufficient) time-ordered directed acyclic graphs (TODAGs) that describe the 3 variables involved in the paradox, several important examples of the paradox need not support this assumption; see sections 2.2.3, 4 and 5. Based on the previous literature, we also argue in section 2 that Simpson’s paradox is sufficiently frequent when the probabilities of the involved variables are generated from the unbiased (non-informative) distribution, modeled via Dirichlet density. Hence this is a genuine decision-making paradox and not an artifact due to inappropriate data gathering.
Our proposal here is to look for the resolution of the paradox by assuming that—given two correlated variables $A$ and $B$—there is a random variable $C$ that makes $A$ and $B$ conditionally independent; i.e., $C$ screens out $A$ from $B$. Examples of Simpson’s paradox show that such a $C$ is frequently plausible, though it is normally not observed directly. In particular, $C$ is conceivable if the correlations between $A$ and $B$ are not caused by a direct causal influence of $B$ on $A$. Then the existence of $C$ is postulated by the common cause principle. (If correlations are caused by a causal influence of $A$ on $B$, Simpson’s paradox can formally exist, but factually it is absent because the decision is obviously to be taken according to the aggregated level.)
Introducing the screening variable $C$ allows us to reformulate and extend Simpson’s paradox: its two options—along with many other options—refer to particular choices of $C$; see section 3. Now the paradox seems to be further from resolution than before. However, we show that when the variables $A_{1}$, $A_{2}$, $B$, and $C$ holding the paradox are binary (the minimal set-up of the paradox), the decision-making is to be made according to the fine-grained probabilities, i.e., the paradox is resolved. Such a definite relation is impossible for a tertiary (or larger) $C$: now depending on $C$ all options of Simpson’s paradox are possible, e.g. the precise control of $C$ can be necessary for decision-making.
Next, we turn to Simpson’s paradox for continuous variables, which was discussed earlier than the discrete formulation [1]. It holds the main message of the discrete formulation. In addition, it includes the concept of the conditional correlation coefficient (only for Gaussian variables is the random-variable dependence fully explained by the correlation coefficient). The continuous formulation is important because it applies to big data [23, 24, 29], and because (statistically) it is more frequent than the discrete version [30]. The advantage of continuous Gaussian formulation is that the general description of the paradox under the common cause is feasible; see section 6. For this situation, we show conceptually the same result as for the discrete version: in the minimal (and most widespread) version of the paradox, the very existence of an (unobservable) common cause leads to preferring the fine-grained option of the paradox.
The rest of this paper is organized as follows. Section 2 is a short but sufficiently inclusive review of Simpson’s paradox and its resolutions proposed in the literature  [^1] Among the issues not addressed in this paper is the explanation of Simpson’s paradox using counterfactual random variables. This subject is reviewed in [6]. . It also discusses two basic examples for illustrating different aspects of the paradox; see section 2.2.3. In section 2.3, we review results about how frequent the paradox is and re-estimate its frequency within an unbiased data generation. In section 3 we reformulate Simpson’s paradox by assuming that there is a common cause (or screening variable) $C$ behind the three variables. Now $C$ need not be observable, since we show that it will be sufficient to assume that it exists and (provided that all variables are binary) Simpson’s paradox is resolved by choosing its fine-grained option. A similar conclusion is reached for Gaussian variables; see section 6. Section 4 considers published data from Ref. [16] on a case of smoking and surviving. This example is not easily treated via the existing methods. Still, we show that the existence of a common cause for this situation is plausible and that Simpson’s paradox can be studied via our method and leads to a reasonable result. Section 5 treats data on COVID-19, which was suggested in Ref. [31]. We demonstrate that an assumption of a plausible common cause points to different conclusions than in Ref. [31]. The last section summarizes and outlines future research directions.
2 Formulation of Simpson’s paradox and previous works
2.1 Formulation of the paradox for binary variables and its necessary conditions
To formulate the paradox in its simplest form, assume three binary random variables $A_{1}=\{a_{1},\bar{a}_{1}\}$, $A_{2}=\{a_{2},\bar{a}_{2}\}$, $B=\{b,\bar{b}\}$. The target event is $a_{1}$, and we would like to know how it is influenced by $A_{2}$ which occurs at an earlier time than the time of $A_{1}$: $t_{A_{2}}\leq t_{A_{1}}$. This can be done by looking at conditional probability. For
$p(a_{1}|a_{2})<p(a_{1}|\bar{a}_{2}),$ (1)
which is equivalent to $p(a_{1})<p(a_{1}|\bar{a}_{2})$, we would conclude that $\bar{a}_{2}$ enables $a_{1}$. However, (1) is compatible with
$p(a_{1}|a_{2},b)$ $>p(a_{1}|\bar{a}_{2},b),$ (2)
$p(a_{1}|a_{2},\bar{b})$ $>p(a_{1}|\bar{a}_{2},\bar{b}),$ (3)
where $B$ also occured in an earlier time: $t_{B}\leq t_{A_{1}}$. Examples supporting (13) are studied below (sections 2.2.3, 4 and 5) and also in Appendix .1. Since (2, 3) hold for each value of $B$ we should perhaps conclude that ${a}_{2}$ enables $a_{1}$ in contrast to (1). Decision-makers would not know whether to apply (1) or (2, 3). This is Simpson’s paradox. Its equivalent formulation is when all inequalities in (13) are inverted  [^2] We leave aside the following pertinent problem; see [19] for details. If probabilities are extracted from finite populations, the more conditioned version (2, 3) is less reliable, because it is extracted from a smaller population. For us all probability-providing populations will be sufficiently large. .
For Simpson’s paradox (13) to hold, it is necessary to have one of the following two conditions:
$p(a_{1}|\bar{a}_{2},b)<p(a_{1}|{a}_{2},b)<p(a_{1}|\bar{a}_{2},\bar{b})<p(a_{1}|{a}_{2},\bar{b}),$ (4)
$p(a_{1}|\bar{a}_{2},\bar{b})<p(a_{1}|{a}_{2},\bar{b})<p(a_{1}|\bar{a}_{2},b)<p(a_{1}|{a}_{2},b).$ (5)
To find these relations, expand $p(a_{1}|a_{2})$ and $p(a_{1}|\bar{a}_{2})$ over the probabilities in (4, 5) [cf. (23, 24)], and note that e.g. $p(a_{1}|a_{2})$ is a weighted mean of $p(a_{1}|{a}_{2},b)$ and $p(a_{1}|{a}_{2},\bar{b})$. Given that (4) or (5) hold, Simpson’s paradox can be generated via suitable choices of $p(b|a_{2})$ and $p(b|\bar{a}_{2})$. For such choices, it is necessary that
$\left(p(b|a_{2})-\frac{1}{2}\right)\left(p(b|\bar{a}_{2})-\frac{1}{2}\right)<0,$ (6)
i.e., $A_{2}$ and $B$ must be dependent variables.
2.2 Attempts to resolve the paradox
2.2.1 Replacing prediction with retrodiction
Over time, several resolutions to the paradox have been proposed. Barigelli and Scozzafava [10, 11] proposed to replace (1) by
$p(a_{2})p(a_{1}|a_{2})<p(\bar{a}_{2})p({a}_{1}|\bar{a}_{2}),$ (7)
i.e. to interchanging $A_{1}$ and $A_{2}$ in (1). Then it is easy to see that its inversion under additional conditioning over $B$ is impossible. While (1) stands for prediction, i.e. aiming at $a_{2}$ (and not at $\bar{a}_{2}$) will more likely produce $\bar{a}_{1}$ (than ${a}_{1}$), the proposal by Ref. [10, 11] looks for retrodiction. Though retrodicting (in contrast to predicting) does not suffer from Simpson’s paradox, retrodicting and predicting are different things, hold different intuitions, and cannot generally be substituted for each other.
Rudas also sought to change the criterion (1) so that it does not allow inversion after additional conditioning over $B$, but still has several reasonable features [32]. The proposal is to employ $p(a_{2})[p(a_{1}|a_{2})-p(\bar{a}_{1}|a_{2})]<p(\bar{a}_{2})[p(a_{1}|\bar{a}_{2})-p(\bar{a}_{1}|\bar{a}_{2})]$ instead of (1) [32]. Notice the conceptual relation of this with the previous proposal (7).
An unnatural point of both these proposals is that they depend on the ratio $p(a_{2})/p(\bar{a}_{2})$; e.g. for the Example 1 mentioned below this means that if the treatment was applied more, it has better chances to be accepted. This drawback is acknowledged in [32].
2.2.2 Exchangeability and causality
According to Lindley and Novick, the paradox may be resolved by going beyond probabilistic considerations (as we do below as well) and by employing the notion of exchangeability or causality [9]. Within that proposal, the data generally provides only propensities, and one needs additional assumptions of sample homogeneity (exchangeability) for equating propensities with probabilities even for a large sample size. Exchangeability and the closely related notion of ergodicity remain influential in the current analysis of statistical problems exemplified by Simpson’s paradox [33]. Lindley and Novick studied the following two examples that support Simpson’s paradox (more examples are discussed in sections 4, 5, and in Appendix .1).
Example 1. Medical treatment [9]. $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the recovery rate of medical patients: $a_{1}={\rm recovery}$, $\bar{a}_{1}={\rm no~{}recovery}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to a specific medical treatment: $a_{2}={\rm treatment}$, $\bar{a}_{2}={\rm no~{}treatment}$. $B=\{b,\bar{b}\}$ is the sex of patients: $b={\rm male}$, $\bar{b}={\rm female}$. The times to which the random variables $A_{1}$, $A_{2}$ and $B$ refer clearly hold $t_{B}<t_{A_{2}}<t_{A_{1}}$.
Example 2. Plant yield [9]. $A_{1}=\{a_{1},\bar{a}_{1}\}$ (the target variable) is the yield of a single plant: $a_{1}={\rm high}$, $\bar{a}_{1}={\rm low}$. $A_{2}=\{a_{2},\bar{a}_{2}\}$ refers to the variety (color) of the plant: $a_{2}={\rm dark}$, $\bar{a}_{2}={\rm light}$. $B=\{b,\bar{b}\}$ refers to the height of the plant: $b_{1}={\rm tall}$, $\bar{b}={\rm low}$. The times hold $t_{A_{2}}<t_{B}<t_{A_{1}}$.
Lindley and Novick proposed that assumptions on exchangeability lead to preferring (1) for Example 2 and (2, 3) for Example 1 [9]. They also proposed that the same results can be found by using causality instead of exchangeability [9]. The same proposal was made earlier by Cartwright in the context of abstract causality [7, 8]. Pearl elaborated this proposal assuming that the above examples can be represented via time-ordered direct acyclic graphs (TODAG) [13, 14], where an arrow $\to$ represents the influence of an earlier variable to the later one; see Fig. 1 for details. If we follow this assumption, then—given the time constraints for the examples—each of them can be related to a unique TODAG:
${\bf Example~{}1}:~{}~{}~{}B\to A_{2}\to A_{1}\leftarrow B,$ (8)
${\bf Example~{}2}:~{}~{}~{}A_{2}\to B\to A_{1}\leftarrow A_{2}.$ (9)
In (8) the suggestion is to condition over $B$ [hence using (2, 3)] if $B$ influences both $A_{1}$ and $A_{2}$ [9, 13, 14]. This is because conditioning over the cause reduces spurious correlations. This reasoning was generalized as the back-door criterion [13].
In contrast, it is advised to use (1) in (9) since $B$ is an effect of $A_{2}$, but still a cause of $A_{1}$ [9, 13, 14]. The intuition of this suggestion is seen in the extreme case when $B$ screens $A_{1}$ and $A_{2}$ from each other, i.e. $A_{1}$, $B$ and $A_{2}$ form a Markov chain. Then the conditional probability $p(A_{1}|A_{2},B)=p(A_{1}|B)$ will not depend on $A_{2}$ begging the original question in (1). Thus, for the two examples considered in [9], Refs. [13, 14] make similar recommendations. The basis of these recommendations was criticized in [17].
2.2.3 Criticism
Realistically, Example 1 need not to support any TODAG. In fact, both arrows $B\to A_{2}$ and $B\to A_{1}$ are generally questionable: sex need not influence the selection of the treatment, $B\not\to A_{2}$ (unless the data was collected in that specific way), and many treatments are sex-indifferent, i.e. $B\not\to A_{1}$. For Example 1 it is more natural to assume that $B$ does not causally influence $A$. In such a situation, the common cause principle proposes that there is an unobserved random variable $C$, which is a common cause for $A$ and $B$ [34, 35]; see section 3.
Similar reservations apply to Example 2: now $A_{2}\to B$ is perhaps argued on the basis of color ($A_{2}$) being more directly related to the genotype of the plant, while the height ($B$) is a phenotypical feature. First, color-genotype and height-phenotype relations need not hold for all plants. Second (and more importantly), it is more natural to assume that the plant genotype influences both its color and height than that the color influences height. Hence the genotype can be a common cause for $A$ and $B$.
2.3 How frequent is Simpson’s paradox: an estimate based on the non-informative Dirichlet density
To estimate the frequency of Simpson’s paradox under fair data-gathering, we can try to generate the probabilities in (13) randomly in an unbiased way, and calculate the frequency of holding the paradox [36, 30]. The best and widely accepted candidate for an unbiased density of probabilities is the Dirichlet density, which is widely employed in statistics and machine learning [37, 38]; see Ref.[39] for a recent review. The Dirichlet probability density for $n$ probabilities $(q_{1},...,q_{n})$ reads:
${\cal D}(q_{1},...,q_{n}|\alpha_{1},...,\alpha_{n})=$
$\frac{\Gamma[\sum_{k=1}^{n}\alpha_{k}]}{\prod_{k=1}^{n}\Gamma[\alpha_{k}]}\prod_{k=1}^{n}q_{k}^{\alpha_{k}-1}\,\delta(\sum_{k=1}^{n}q_{k}-1),$ (10)
$\int_{0}^{\infty}\prod_{k=1}^{n}{\rm d}q_{k}{\cal D}(q_{1},...,q_{n}|\alpha_{1},...,\alpha_{n})=1,$ (11)
where $\alpha_{k}>0$ are the parameters of the Dirichlet density, $\delta(x)$ is the delta-function, and $\Gamma[x]=\int_{0}^{\infty}{\rm d}q\,q^{x-1}e^{-q}$ is the Euler’s $\Gamma$-function. Since ${\cal D}(q_{1},...,q_{n})$ is non-zero only for $q_{k}\geq 0$ and $\sum_{k=1}^{n}q_{k}=1$, the continuous variables themselves have the meaning of probabilities.
Many standard prior densities for probabilities are contained in (2.3); e.g., homogeneous ($\alpha_{1}=...\alpha_{n}=1$), Haldane’s ($\alpha_{1}=...\alpha_{n}=\alpha\approx 0$), Jeffreys ($\alpha_{1}=...\alpha_{n}=1/2$). For estimating the frequency of Simpson’s paradox, Ref. [36] employed homogeneous and Jeffreys prior.
For modeling a non-informative Dirichlet density we find it natural to take
$\alpha_{1}=...\alpha_{n}={1}/{n}.$ (12)
The homogeneity feature, $\alpha_{1}=...\alpha_{n}$ in (12) is natural for an unbiased density. The factor $\frac{1}{n}$ in (12) makes an intuitive sense, since $\alpha_{1}=...\alpha_{n}$ become homogeneous (non-informative) probabilities. Eq. (12) arises when we assume that the distribution of random probabilities is independent of whether they were generated directly from (2.3) with $n$ components, or alternatively from (2.3) with $nm$ components $\alpha_{1}=....=\alpha_{nm}$, and then marginalized. This requirement indeed leads to (12), as can be checked with the following feature of (2.3):
$\int_{0}^{\infty}{\rm d}q_{n-1}^{\prime}\,{\rm d}q_{n}^{\prime}\,\delta(q_{n-1}-q_{n-1}^{\prime}-q_{n}^{\prime})\,\times$
${\cal D}(q_{1},...,q_{n-2},q_{n-1}^{\prime},q_{n}^{\prime}|\alpha_{1},...,\alpha_{n})$
$={\cal D}(q_{1},...,q_{n-2},q_{n-1}|\alpha_{1},...,\alpha_{n-2},\alpha_{n-1}+\alpha_{n}).$ (13)
The message of (13) is that aggregating over two probabilities leads to the same Dirichlet density with the sum of the corresponding weights $\alpha_{n-1}$ and $\alpha_{n}$.
We estimated the frequency of Simpson’s paradox assuming that 8 probabilities $p(A_{1},A_{2},B)$ in (13) are generated from (2.3, 12) with $n=8$ (binary situation). This amounts to checking two relations (they amount to (13) and its reversal)
$[p(a_{1}|a_{2})-p(a_{1}|\bar{a}_{2})][p(a_{1}|a_{2},b)-p(a_{1}|\bar{a}_{2},b)]<0,$ (14)
$[p(a_{1}|a_{2},b)-p(a_{1}|\bar{a}_{2},b)][p(a_{1}|a_{2},\bar{b})-p(a_{1}|\bar{a}_{2},\bar{b})]>0.$
Our numerical result is that the frequency of two inequalities in (14) is $\approx 4.29\%\pm 0.001\%$. For this precision it was sufficient to generate $N=10^{7}$ samples from (2.3, 12) with $n=8$. This result compares favorably with $\approx 1.66\%$ obtained for $\alpha_{1}=...\alpha_{8}=1$ (homogeneous prior), and $\approx 2.67\%$ obtained for $\alpha_{1}=...\alpha_{8}=0.5$ (Jeffreys prior) [36]. It is seen that the frequency of Simpson’s paradox is a decreasing function of $\alpha_{1}=...=\alpha_{8}=\alpha$ [36].
Roughly, the above result $\approx 4.29\%$ means that in every 1000 instances of 3 binary variables, 42 instances will show Simpson’s paradox. This number is reassuring: it is not very large meaning that the standard decision-making based on the marginal probabilities in (1) will frequently be reasonable. But it is also not very small, showing that Simpson’s paradox is generic and has its range of applicability.
3 Common cause principle and reformulation of Simpson’s paradox
Figure 1: Directed acyclic graphs between random variables $A=(A_{1},A_{2})$, $B$ and $C$ involved in discussing Simpson’s paradox. The first and second graphs were studied in Refs. [13, 14]; see (8, 9). The third or fourth graphs are basic assumptions of this work; see (15). In the first graph, $B$ influences $A_{1}$ and $A_{2}$, but $B$ is not the common cause in the strict sense, because there is an influence from $A_{2}$ to $A_{1}$. A similar interpretation applies to the second graph. We emphasize that the joint probability $p(A_{1},A_{2},B)$ for the first and second graphs has the same form, i.e. such graphs are extra constructions employed for interpretation of data. In contrast, the third and fourth graph imply a definite (but the same for both graphs) limitation on the joint probability $p(A_{1},A_{2},B,C)$, which is expressed by (15).
3.1 Common cause and screening
The common cause for $A=(A_{1},A_{2})$ and $B$ means that there exists a random variable $C=\{c\}$ [34, 35]
$p(A_{1},A_{2},B|C)=p(A_{1},A_{2}|C)p(B|C),$ (15)
$p(A_{1},A_{2},B)={\sum}_{c}\,p(A_{1},A_{2}|c)p(B|c),$ (16)
where (15) holds for all values assumed by $A_{1}$, $A_{2}$, $B$ and $C$, and where (16) follows from (15)  [^3] There are formulations of the common cause principle that look for (15) holding for certain events only and not for random variables [34, 35]. We do not focus on them. . The same (15) applies if $C$ causes $A$ and screens $A$ from $B$. These two scenarios are shown in Fig. 1 as (resp.) the third and fourth graphs. Sections 4, 5 and Appendix .1 provide several examples of a causing (or screening) variable $C$ in the context of Simpson’s paradox. The common cause principle was proposed to explain probabilistic correlations [34, 35]. It later found important applications in data science, where approximate relations similar to (15) are applied to effective data compression (Non-negative matrix factorization, Probabilistic Latent Dirichlet indexing, etc); see [40] for a review.
Note from (15) that $C$ gets rid of the conditional dependence on $B$ in $p(A_{1},A_{2}|B,C)$. Thus, a sensible way of looking at the association between $a_{1}$ and $a_{2}$ is to check the sign of
$p(a_{1}|a_{2},C)-p(a_{1}|\bar{a}_{2},C)~{}~{}~{}~{}{\rm for~{}each~{}value~{}of~{}}C.$ (17)
To support the usage of the common cause C for decision-making, we note that (15) has an important implication in the context of (1). (This implication generalizes the argument given in [35].) Assume that $p(a_{2},b,c)>0$ for all values $c$ of $C$. Note from (15) that there exists an event $c$ such that $p(a_{1}|a_{2},b)\leq p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$, and an event $c^{\prime}$ such that $p(a_{1}|a_{2},b)\geq p(a_{1}|a_{2},b,c^{\prime})=p(a_{1}|a_{2},c^{\prime})$. Hence, if conditioning over $b$ facilitates (hinders) the association between $a_{1}$ and $a_{2}$, then conditioning over $c$ ($c^{\prime}$) is not worse in this facilitation (hindering)  [^4] To deduce the first relation assume that $p(a_{1}|a_{2},b)<p(a_{1}|a_{2},b,c)=p(a_{1}|a_{2},c)$ for all $c$, multiply both parts by $p(a_{2},b,c)>0$, sum over $c$ and get contradiction $p(a_{1},a_{2},b)<p(a_{1},a_{2},b)$. Likewise for the second relation..
After the above reformulation, Simpson’s paradox seems even less resolvable since $C$ is not observed. Indeed, there are common causes that reproduce (1), those that reproduce (2, 3), but there are many other possibilities. Common causes that are close to $B$ ($C\approx B$) imply option (2, 3) of the paradox, while $C\approx A$ leads to (1). These conclusions are based on the fact that (15) holds exactly for $C=B$ and $C=A$. Thus, Simpson’s paradox is not a choice between two options (2, 3) and (1), it is a choice between many options given by different common causes $C$.
Finally, two remarks about the applicability of (1517). First, if $C$ is a common cause for both $A=(A_{1},A_{2})$ and $B$, the times of these variables naturally hold $t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}},t_{B}]$. When $C$ screens $A$ from $B$, it holds $t_{B}<t_{C}<{\rm min}[t_{A_{1}},t_{A_{2}}]$. In certain applications of (17), it will suffice to have even a weaker condition $t_{C}<t_{A_{1}}$.
Second, we note that for applying (1, 2, 3) we do not need $p(A_{2})$, i.e. only $p(B|A_{2})$ is needed for connecting (1) with (2, 3). Indeed, $A_{2}$ does not necessarily need to be a random variable, but can simply be a label describing the situation. Now the same holds for (17): once (15) is written as
$p(A_{1},B|A_{2},C)=p(A_{1}|A_{2},C)p(B|C),$ (18)
we need only $p(C|A_{2})$ to pass from (18) to quantities involved in (1, 2, 3); i.e., $p(A_{2})$ is not needed.
3.2 A common cause (or screening variable) resolves Simpson’s paradox for binary variables
The following theorem shows a definite statement for all binary causes. The message of the theorem is that once we know that $C$ is binary, then the correct decision is (2, 3).
Theorem 1: If $A_{1}$, $A_{2}$, $B$ and $C=\{c,\bar{c}\}$ are binary, and provided that (1) and (2, 3) are valid, all causes $C$ hold
$p(a_{1}|a_{2},c)>p(a_{1}|\bar{a}_{2},c),\qquad p(a_{1}|a_{2},\bar{c})>p(a_{1}|\bar{a}_{2},\bar{c}),$ (19)
i.e. all $C$ holding (15) predict the same sign of association between $a_{1}$ and $a_{2}$ as (2, 3).
The main idea of proving (19) is inverting (15):
$p(a_{1}|a_{2},c)$
$=\frac{p(\bar{b}|\bar{c})p(a_{1}|a_{2},b)p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2})}{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}$ (20)
$=p(a_{1}|a_{2},\bar{b})+\frac{p(\bar{b}|\bar{c})p(b|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]}{1-p(\bar{b}|\bar{c})-p(b|a_{2})},$ (21)
$p(c|a_{2})=\frac{p(\bar{b}|\bar{c})p(b|a_{2})+(p(\bar{b}|\bar{c})-1)p(\bar{b}|a_{2})}{p(b|c)+p(\bar{b}|\bar{c})-1}$
$=\frac{p(\bar{b}|\bar{c})+p(b|a_{2})-1}{p(b|c)+p(\bar{b}|\bar{c})-1},$ (22)
where unknown quantities $p(a_{1}|a_{2},c)$ and $p(c|a_{2})$ are represented via known ones (i.e. $p(A_{1},A_{2},B)$) and free parameters $p(B|C)$. Eqs. (21, 22) hold upon changing $a_{2}$ by $\bar{a}_{2}$ and are deduced in Appendix .3 via specific notations that should be useful when dealing with (15) for a non-binary $C$.
The rest of the proof is algebraic but non-trivial. It also works out and employs constraints (4, 31) on Simpson’s paradox itself. Expanding both sides of (1),
$p(a_{1}|a_{2})=p(a_{1}|a_{2},b)p(b|a_{2})+p(a_{1}|a_{2},\bar{b})p(\bar{b}|a_{2}),$ (23)
$p(a_{1}|\bar{a}_{2})=p(a_{1}|\bar{a}_{2},b)p(b|\bar{a}_{2})+p(a_{1}|\bar{a}_{2},\bar{b})p(\bar{b}|\bar{a}_{2}),$ (24)
and using there (2, 3) we subtract the sides of (1) from each other and find:
$p(a_{1}|a_{2},b)+p(\bar{b}|a_{2})[p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b)]<$
$p(a_{1}|\bar{a}_{2},b)+p(\bar{b}|\bar{a}_{2})[p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)].$ (25)
We return to (2, 3) and note that we can assume without loosing generality
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|a_{2},b).$ (26)
Eqs. (23, 24) imply that for the validity of (13, 26) it is necessary to have $p(a_{1}|\bar{a}_{2},\bar{b})>p(a_{1}|a_{2},b)$, which together with (2, 3, 26) revert to (4). Now (1, 23, 24) read
$p(a_{1}|a_{2},b)+[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))$ (27)
$<p(a_{1}|\bar{a}_{2},b)+[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$
$p(a_{1}|a_{2},\bar{b})-p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))$
$<p(a_{1}|\bar{a}_{2},\bar{b})-p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (28)
where (27) and (28) are equivalent. Eqs. (27, 28, 4) imply
$[1-p(b|a_{2})](p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))<$
$[1-p(b|\bar{a}_{2})](p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)),$ (29)
$p(b|a_{2})(p(a_{1}|a_{2},\bar{b})-p(a_{1}|a_{2},b))>$
$p(b|\bar{a}_{2})(p(a_{1}|\bar{a}_{2},\bar{b})-p(a_{1}|\bar{a}_{2},b)).$ (30)
As checked directly, Eqs. (3.2, 30) lead to
$p(b|a_{2})>p(b|\bar{a}_{2}).$ (31)
Now we return to (22) and assume there ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$, which leads to ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}<0$ from (22). Writing down from (22) the formula for $p(c|\bar{a}_{2})$ and making the same assumption we get ${p(\bar{b}|\bar{c})+p(b|\bar{a}_{2})-1}<0$. Now look at (21) and its analog obtained via $a_{2}\to\bar{a}_{2}$, and use there these two results together with (30, 31) and (4) to deduce the first inequality in (19) under assumption ${p(b|c)+p(\bar{b}|\bar{c})-1}<0$. It should be obvious that the second inequality in (19) holds under the same assumption since we nowhere used any specific feature of $c$ compared to $\bar{c}$.
For ${p(\bar{b}|\bar{c})+p(b|a_{2})-1}>0$ we need to use instead of (21) another form of (20)
$p(a_{1}|a_{2},c)=p(a_{1}|a_{2},b)-$
$\frac{[1-p(\bar{b}|\bar{c})]p(\bar{b}|a_{2})[p(a_{1}|a_{2},b)-p(a_{1}|a_{2},\bar{b})]}{p(\bar{b}|\bar{c})+p(b|a_{2})-1}.$ (32)
The rest is similar to the above: we proceed via (3.2, 31) and (4) and deduce (19) from (22), (3.2) and the analog of (3.2) obtained via $a_{1}\to\bar{a}_{2}$.
3.3 Non-binary causes
Starting from a tertiary $C$, all three options of Simpson’s paradox become possible: there are common causes $C$ that support (1), those which support (2, 3), and eventually (random) cause variables $C=\{c_{1},c_{2},c_{3}\}$ for which $p(a_{1}|a_{2},c_{i})-p(a_{1}|\bar{a}_{2},c_{i})$ has different signs for different values of $i=1,2,3$. (Our numerical examples showing these possibilities are available upon request.) Hence, already for the tertiary cause one needs prior information on the common cause to decide on the solution of Simpson’s paradox. Alternatively, we can infer this unknown cause via e.g. one of the methods proposed recently [41, 42]. It is not excluded that such inference methods will provide further information on the solution of Simpson’s paradox. We hope to discuss this problem elsewhere.
4 Example: smoking and surviving
In sections 2.2.2 and 2.2.3 we discussed two examples studied in the literature and argued that they can be also interpreted via the common cause principle. In the present case, the standard approaches do not seem to apply, but the common cause can still be motivated. This example on survival of smokers versus nonsmokers is taken from Ref. [16]. Its technical details are discussed in Appendix .2.
Binary $A_{1}$ represents the survival in a group of women as determined by two surveys taken 20 years apart:
$A_{1}=\{a_{1},\bar{a}_{1}\}=\{{\rm died},{\rm alive}\},$
$A_{2}=\{a_{2},\bar{a}_{2}\}=\{{\rm smoker},{\rm nonsmoker}\},$ (33)
[…]
where $p(\bar{b})=0.1334$, and where $b$ and $\bar{b}$ denote age-groups. According to the data of [16], Simpson’s paradox reads
$p(a_{1}|a_{2})=0.2214<p(a_{1}|\bar{a}_{2})=0.2485,$ (35)
$p(a_{1}|a_{2},b)>p(a_{1}|\bar{a}_{2},b),$
$p(a_{1}|a_{2},\bar{b})>p(a_{1}|\bar{a}_{2},\bar{b}).$ (36)
Note that $B$ here influences $A_{1}$: the age of a person is a predictor of his/her survival. Causal influences from age to smoking can be neglected because the number of people that quit or started smoking is small [16]. We can assume that influences from smoking to age are absent. Then this example is intermediate between two situations considered in [7, 8, 9, 13]. Recall that when $B$ influenced $A_{2}$, these references advised to decide via the fine-grained option of the paradox, while for the case of the inverse influence (from $A_{2}$ to $B$) they recommend to employ the coarse-grained version; see Fig. 1.
Hence, we should expand on the above situation to achieve a workable model. We can assume that $A_{2}$ and $B$ are influenced by a common cause. Genetic factors influence an individual’s age and tendency to smoke. Originally proposed by Fisher [43], this hypothesis was later substantiated in several studies; see Refs. [44, 45] for reviews. Note that this refers to genetics of the smoking behavior itself, and not to health problems that can be caused by smoking plus genetic factors. Several sets of studies that contributed to genetic determinants of smoking behavior are as follows. (i) Children of smoking parents tend to smoke. (ii) Smoking behavior of adopted kids correlates stronger with that of their biological parents. (iii) Monozygotic (genetically identical) twins correlate in their smoking behavior much stronger than heterozygotic twins. Smoking behavior includes both the acquisition and maintenance of smoking. Monozygotic twins show correlations in both these aspects.
Hence as a preliminary hypothesis, we can suggest that genetic factors are the common cause of both smoking and age. If this common cause is binary, then Theorem 1 applies and we conclude—judging from the fine-grained data and consistently with other studies—that smoking is not beneficial for surviving.
5 Example: COVID-19, Italy versus China
Here the COVID-19 death rates are compared in Italy and China [31, 46]. According to the data, aggregated death rates in Italy are higher than in China, but in each age group, the death rates are higher in China. More precisely,
$A_{1}=\{a_{1},\bar{a}_{1}\}=\{{\rm died},{\rm alive}\},$
$A_{2}=\{a_{2},\bar{a}_{2}\}=\{{\rm China},{\rm Italy}\},$ (37)
[…]
where $p(a_{1})$ is the death rate out of COVID-19, $p(B)$ is found from the number of positively tested people in each age group, $p(\bar{b})=0.1012$, and where $p(\bar{b}|a_{2})=0.1017$ and $p(\bar{b}|\bar{a}_{2})=0.3141$. According to the data of [31], Simpson’s paradox reads
$p(a_{1}|a_{2})=0.0608<p(a_{1}|\bar{a}_{2})=0.0760,$ (39)
$p(a_{1}|a_{2},b)=0.0507>p(a_{1}|\bar{a}_{2},b)=0.04900,$
$p(a_{1}|a_{2},\bar{b})=0.150>p(a_{1}|\bar{a}_{2},\bar{b})=0.135.$ (40)
The authors of [31] proposed that this situation is described by TODAG $A_{2}\to B\to A_{1}\leftarrow A_{2}$; cf. (9). Then the conclusion from [9, 13] will be that the aggregated version of Simpson’s paradox works, i.e. Italy did worse than China. The authors of Ref. [31] reached the same conclusion.
When applying the common cause set-up from section 3.1, we can look at (18), because $A_{2}$ is better described as a label (avoiding dealing with the probability of country). Hence, from the viewpoint of (18), we need a common cause that supplements $A_{2}$ and acts on both $A_{1}$ and $B$. We propose that the quality of healthcare system can be the common cause $C$ here. In particular, a more affordable healthcare system may cause a higher proportion of older people in the country’s society. Indeed, for 2019, Italy had a larger percentage of people aged above 65 than China: 24.05 % versus 12.06 %. On the other hand, the healthcare system will influence death rates in all age groups. If $C$ is binary, then our conclusion from Theorem 1 is opposite to that of [31]: China did worse than Italy.
6 Simpson’s paradox and common cause principle for Gaussian variables
6.1 Formulation of Simpson’s paradox for continuous variables
Simpson’s paradox is uncovered earlier for continuous variables than for the discrete case [1]. Researching the continuous variable paradox and identifying it in big datasets is currently an active research field [23, 24, 29, 47, 48, 49].
The association between continuous variables $A_{1}=\{a_{1}\}$ and $A_{2}=\{a_{2}\}$ can be based on a reasonable definition of correlation coefficient [1, 30]. We focus on Gaussian variables, because this definition is unique for them and amounts to conditional variance. These variables are also important in the context of machine learning (e.g. linear regressions) [50].
Hence the formulation of Simpson’s paradox given $B=\{b\}$ reads instead of (13) [1, 30, 23, 24]:
$\sigma[a_{1},a_{2}]\sigma[a_{1},a_{2}|b]<0~{}~{}{\rm for~{}all~{}}b,$ (41)
$\sigma[a_{1},a_{2}]\equiv\langle(a_{1}-\langle a_{1}\rangle)(a_{2}-\langle a_{2}\rangle)\rangle,$ (42)
$\sigma[a_{1},a_{2}|b]\equiv\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b},$ (43)
$\langle a\rangle_{b}\equiv\int{\rm d}a\,a\,p(a|b),$ (44)
where $\langle a\rangle_{b}$ and $\sigma[a_{1},a_{2}|b]$ are the conditional mean and covariance; $\langle a\rangle$ and $\sigma[a_{1},a_{2}]$ are the mean and covariance; $p(a|b)$ is the conditional probability density of $A=\{a\}$.
The message of (41) is that the usual and conditional covariance have different signs, i.e., they predict different types of associations between $A_{1}$ and $A_{2}$. For instance, $\sigma[a_{1},a_{2}]>0$ means correlation, while $\sigma[a_{1},a_{2}|b]$ implies anti-correlation. Note a subtle difference between this formulation of Simpson’s paradox and that presented in section 2.2. In (4142) the formulation is symmetric with respect to $A_{1}$ and $A_{2}$.
6.2 General solution for Gaussian variables
For fuller generality, we shall assume that $A=\{\boldsymbol{a}\}$, and $B=\{\boldsymbol{b}\}$ are Gaussian column vectors with a number of components (i.e., dimensionality) $n_{A}$, and $n_{B}$, respectively. We also define
$\boldsymbol{y}^{\rm T}=(\boldsymbol{a},\boldsymbol{b}),$ (45)
where ${\rm T}$ means transposition: $\boldsymbol{y}^{\rm T}\boldsymbol{y}$ is a number, while $\boldsymbol{y}\boldsymbol{y}^{\rm T}$ is a matrix. We assume that a Gaussian $n_{X}$-dimensional variable $X=\{\boldsymbol{x}\}$ is the common cause variable for $A$ and $B$:
$P(\boldsymbol{y}|\boldsymbol{x})=\frac{1}{(2\pi)^{n_{A}/2}\,\sqrt{{\rm det}[{\cal Q}]}}\,\,e^{-\frac{1}{2}(\boldsymbol{y}-{\cal C}\boldsymbol{x})^{\rm T}{\cal Q}^{-1}(\boldsymbol{y}-{\cal C}\boldsymbol{x})},$
$P(\boldsymbol{x})=\frac{1}{(2\pi)^{n_{X}/2}\sqrt{{\rm det}[{\cal S}]}}\,\,e^{-\frac{1}{2}\boldsymbol{x}^{\rm T}{\cal S}^{-1}\boldsymbol{x}},$ (47)
${\cal Q}=\begin{pmatrix}{\cal A}&0\\ 0&{\cal B}\end{pmatrix},\qquad\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}={\cal C}\boldsymbol{x},$ (48)
$\langle(\boldsymbol{y}-\langle\boldsymbol{y}\rangle_{\boldsymbol{x}})(\boldsymbol{y}^{\rm T}-\langle\boldsymbol{y}^{\rm T}\rangle_{\boldsymbol{x}})\rangle_{\boldsymbol{x}}={\cal Q},$ (49)
where the common cause feature of $X=\{\boldsymbol{x}\}$ is ensured by the block-diagonal structure of the covariance matrix ${\cal Q}$: ${\cal A}$ and ${\cal B}$ are (resp.) covariance matrices for $A$ and $B$. In (6.2), ${\cal C}$ is $(n_{A}+n_{B})\times n_{X}$ matrix that ensures the coupling between $(A,B)$ and $X$. For simplicity and without loss of generality we assumed that $\langle\boldsymbol{x}\rangle=0$ and hence $\langle\boldsymbol{y}\rangle=0$ in (6.2). We get from (6.2) after arranging similar terms (and omitting normalization):
$P(\boldsymbol{x})P(\boldsymbol{y}|\boldsymbol{x})\propto\exp\bigg{(}-\frac{1}{2}[\,\boldsymbol{x}^{\rm T}-\boldsymbol{y}^{\rm T}{\cal Q}^{-1}{\cal C}V^{-1}\,]$
×$\times\,V\,[\,\boldsymbol{x}-V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}\boldsymbol{y}]$
$-\frac{1}{2}\boldsymbol{y}^{\rm T}[\,{\cal Q}^{-1}-{\cal Q}^{-1}{\cal C}V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}\,]\boldsymbol{y}\,\bigg{)},$ (50)
$V={\cal S}^{-1}+{\cal C}^{\rm T}{\cal Q}^{-1}{\cal C}.$ (51)
Employing (84) we obtain:
${\cal Q}^{-1}-{\cal Q}^{-1}{\cal C}V^{-1}{\cal C}^{\rm T}{\cal Q}^{-1}=({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1},$ (52)
$P(\boldsymbol{y})\propto e^{-\frac{1}{2}\boldsymbol{y}^{\rm T}\,({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1}\,\boldsymbol{y}\,},$ (53)
$\langle\boldsymbol{y}\boldsymbol{y}^{\rm T}\rangle={\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T},$ (54)
We now recall (45, 49), introduce the block-diagonal form for ${\cal C}{\cal S}{\cal C}^{\rm T}$, and find
${\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T}=\begin{pmatrix}{\cal A}+{\cal J}&{\cal K}\\ {\cal K}^{\rm T}&{\cal B}+{\cal L}\end{pmatrix},$ (55)
$({\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T})^{-1}=\begin{pmatrix}({\cal A}+{\cal J}-{\cal K}({\cal B}+{\cal L})^{-1}{\cal K}^{\rm T})^{-1}&...\\ ...&...\end{pmatrix},$
where (LABEL:woot) can be deduced via Appendix .4. In that formula we need only the upper-left block, so that all other blocks are omitted. Collecting pertinent expressions from (45, 54, LABEL:woot, 49), we deduce finally
$\langle\boldsymbol{y}\boldsymbol{y}^{\rm T}\rangle={\cal Q}+{\cal C}{\cal S}{\cal C}^{\rm T},$ (57)
$\langle(\boldsymbol{y}-\langle\boldsymbol{y}\rangle_{\boldsymbol{x}})(\boldsymbol{y}^{\rm T}-\langle\boldsymbol{y}^{\rm T}\rangle_{\boldsymbol{x}})\rangle_{\boldsymbol{x}}={\cal Q},$ (58)
$\langle(\boldsymbol{a}-\langle\boldsymbol{a}\rangle_{\boldsymbol{b}})(\boldsymbol{a}^{\rm T}-\langle\boldsymbol{a}^{\rm T}\rangle_{\boldsymbol{b}})\rangle_{\boldsymbol{b}}$
$={\cal A}+{\cal J}-{\cal K}({\cal B}+{\cal L})^{-1}{\cal K}^{\rm T}.$ (59)
6.3 The minimal set-up of Simpson’s paradox: 3 scalar variables + scalar cause
For this simplest situation, $\boldsymbol{y}^{\rm T}=(a_{1},a_{2},b)$ is a 3-dimensional vector, ${\cal A}$ is a $2\times 2$ matrix, ${\cal C}$ is a $3\times 1$ matrix, while ${\cal S}$ and ${\cal B}$ are positive scalars. Now (5759) read:
$\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon,$
$0<\epsilon\equiv\frac{{\cal B}}{{\cal B}+{\cal C}^{2}_{31}{\cal S}}<1,$ (60)
$\langle(a_{1}-\langle a_{1}\rangle_{x})(a_{2}-\langle a_{2}\rangle_{x})\rangle_{x}={\cal A}_{12},$ (61)
$\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}.$ (62)
Now consider a scenario of Simpson’s paradox, where
$\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}>0~{}~{}{\rm and}~{}~{}$ (63)
$\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon<0.$ (64)
Due to $10<\epsilon<$ 1, these two inequalities demand ${\cal A}_{12}<0$. Likewise, $\langle a_{1}a_{2}\rangle={\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}<0$ and ${\cal A}_{12}+{\cal C}_{11}{\cal C}_{21}{\cal S}\,\epsilon>0$ demand ${\cal A}_{12}>0$. It is seen that under Simpson’s paradox for this minimal situation, the sign of $\langle(a_{1}-\langle a_{1}\rangle_{x})(a_{2}-\langle a_{2}\rangle_{x})\rangle_{x}$ coincides with the sign of $\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}$. We are thus led to the following:
Theorem 2: In the minimal situation (6.362) with the (minimal) common cause, the continuous Simpson’s paradox (41) is resolved in the sense that the decision on the sign of correlations should proceed according to the fine-grained option: $\langle(a_{1}-\langle a_{1}\rangle_{b})(a_{2}-\langle a_{2}\rangle_{b})\rangle_{b}$; see (4142).
For non-minimal common causes, all possibilities of the paradox can be realized; see Appendix .5.
7 Conclusion
We addressed Simpson’s paradox: the problem of setting up an association between two events $a_{1}$, $a_{2}$ given the lurking variable $B$. This decision-making paradox provides two plausible but opposite suggestions for the same situation; see (1) and (2, 3). Either the first option is correct, the second option is correct, or none of them is correct.
We focus on cases when there is a common cause $C$ for $B$ and $A=(A_{1},A_{2})$ (which combines $a_{1}$, $a_{2}$ and their complements). Alternatively, $C$ screens out $A$ from $B$. These cases include those in which there is no causal influence from $A$ to $B$, as well as from $B$ to $A$. Hence, correlations between $A$ and $B$ are to be explained via the common cause $C$, which is a statement of the common cause principle [34, 35]. Now the association between $a_{1}$ and $a_{2}$ is to be decided by looking at $p(a_{1}|a_{2},c)$ for various values of $C$. This task is normally difficult given the fact that $C$ is frequently not fully known and is not observed. However, provided that $A_{1}$, $A_{2}$, $B$ and $C$ are binary, $p(a_{1}|a_{2},c)$ shows the same association as the option (2, 3) of Simpson’s paradox. In this sense, Simpson’s paradox is resolved in the binary situation, provided that the situation allows a binary cause or a binary screening variable. The same conclusion on resolving Simpson’s paradox was reached for Gaussian variables in the minimal situation. Several examples can illustrate the plausibility of a minimal $C$.
These results lead to several interesting research directions. First, in the present paper, we limited ourselves to results that hold for all (minimal) common causes. For many applications this is too stringent: if the common cause is known to exist, but is not observed directly, then it may be sufficient to infer it e.g. via the (generalized) maximum likelihood [42] or the minimal entropy method [41]. This may provide pertinent information on the real common cause and on the structure of Simpson’s paradox. Second, we insisted on a precise common cause. The screening relation (16) is also useful, when it does hold approximately, but the support of $C$ is relatively small. Such an approximate relation (16) provides data compression via feature detection, which is the main message of unsupervised methods such as Non-negative Matrix factorization and Probabilistic Latent Dirichlet indexing [40]. The impact of such approximate, but efficient causes on probabilistic reasoning is an interesting research subject that we plan to explore in the future. Third, the general formalism we developed in section 6 for Gaussian variables may find further applications in the causal analysis of Gaussian machine learning algorithms [50].